OpenStudy (anonymous):

(n+6)(n+7) would the product of this be n^2+n+42??

OpenStudy (anonymous):

Nope. The product of the inner and outer terms don't add to be n.

OpenStudy (anonymous):

ohh you add the inner terms instead of subtracting them

OpenStudy (anonymous):

(a+b)(c+d) = (a+b)c + (a+b)d = ac + ad + bc + bd

OpenStudy (anonymous):

Use the FOIL method (First, Outer, Inner, Last). You multiply the two n's first (Your first terms), Then you multiply n & 7 (Your outer terms). Then you multiply n & 6 (Your inner terms) and then you multiply 6 & 7 (Your last terms)

OpenStudy (anonymous):

You have to sum the products you get from the foil method. Sometimes it ends up being the sum of a negative number which leads to a middle term that's very small (or zero) but you're always adding these products together.